OpenStudy (anonymous):

Help. Explain how you would use the substitution method to solve the following system of equations. 5x - y = -2 10x - 3y = -7

OpenStudy (anonymous):

Start with the top equation. Express it in terms of y. \[5x-y=-2\] \[-y=-5x-2\] \[y=5x+2\]

OpenStudy (anonymous):

Now we can substitute the solved value of y of the top equation to the bottom equation. 10x-3y=-7 10x-3(5x+2)=-7

OpenStudy (anonymous):

Then solve the equation for x. 10x-3(5x+2)=-7 10x-15x-6=-7 -5x-6=-7 -5x=-1 x=1/5 You can solve for y by substituting x = 1/5 to any of the original equations. For more information: http://chilimath.com/algebra/intermediate/subs/substitution-method.html